# LLM4DSR: Leveraing Large Language Model for Denoising Sequential Recommendation

### v1

@@ 5.1. Sequential Recommendation
The sequential recommendation infers the next item of interest to the user based on their historical interactions. Compared to collaborative filtering, sequential recommendation takes into account the temporal order of interactions to better capture the dynamic changes in user interests. Benefiting from the development of deep learning models in recent years, sequential recommendation employs various deep models to model users’ historical interactions and capture their interests. For instance, GRU4Rec (Hidasi et al., 2015) uses RNN, while Caser (Tang and Wang, 2018) uses CNN. SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) are based on the self-attention mechanism (Vaswani, 2017), which automatically learns the weight of each interaction. DROS (Yang et al., 2023) introduces distributionally robust optimization (DRO) (Rahimian and Mehrotra, 2019) to enhance the model’s ability to handle out-of-distribution (OOD) problems. The readers may refer to the excellent survey (Fang et al., 2020; Wang et al., 2019) for more details. However, the aforementioned methods do not consider the negative impact of noisy interactions on modeling the sequence, which is quite common in real-world scenarios.
@@ 5.2. Denoising Recommendation
Regardless of whether explicit or implicit denoising methods are used, the most significant challenge stems from the lack of effective supervision signals from noise labels to train the denoising model. Existing methods often rely on heuristic designs inspired by prior knowledge to develop denoising algorithms, which requires substantial expert knowledge and lacks precision. Additionally, since these methods usually utilize supervision information from recommendation loss, they are still adversely affected by noise as the labels of the recommendation label (i.e., the next item) could also be contaminated. To address these issues, LLM4DSR leverages the extensive open knowledge of Large Language Models (LLMs) to aid in more effective denoising.
Additionally, some works focus on denoising under collaborative filtering scenarios (Wang et al., 2021a, c; Gao et al., 2022; Tian et al., 2022; Wang et al., 2022; Fan et al., 2023; Wang et al., 2023a; He et al., 2024; Li et al., 2024c). Compared to denoising in sequential recommendation, these methods do not take the temporal information of interactions into account. Readers can refer to the survey (Zhang et al., 2023a) for more details.
@@ 5.3. LLMs for Recommendation
The initial attempts directly use pre-trained LLMs as the backbone for recommendations to explore their zero-shot capabilities. (Gao et al., 2023; Hou et al., 2024; Liu et al., 2024; Wang and Lim, 2023; Wang et al., 2024b). These methods perform not well due to the significant gap between recommendation tasks and the training tasks of LLMs. Subsequent research has organized recommendation data into prompt formats and then fine-tuned LLMs to enhance their recommendation performance, achieving better results (Bao et al., 2023a, b; Hu et al., 2024; Li et al., 2023a, 2024b, b, 2024a; Shi et al., 2024; Wang et al., 2024a; Zhang et al., 2023b, 2024b). Besides generating recommendation results directly, LLMs can also serve as enhancers to improve the performance of traditional recommendation models (Xi et al., 2023; Ren et al., 2024; Wei et al., 2024; Yang et al., 2024; Wang et al., 2023b).
To the best of our knowledge, this paper is the first to explore the application of LLMs in the field of sequence recommendation denoising.
@@ 6. Conclusion
Ethical Considerations
Denoising methods essentially transfer the model’s prior knowledge to the data. If the model’s outputs are biased, for example, if it tends to recommend items of a specific category, it will further amplify the biases present in the data. In the context of using LLMs for denoising as discussed in this paper, it is crucial to ensure that the outputs of the pre-trained LLM are inherently fair. Special attention must be paid to bias testing when selecting the backbone model.

### v2

@@ 5.1. Sequential Recommendation
The sequential recommendation infers the next item of interest to the user based on their historical interactions. Compared to collaborative filtering, sequential recommendation takes into account the temporal order of interactions to better capture the dynamic changes in user interests. Benefiting from the development of deep learning models in recent years, sequential recommendation employs various deep models to model users’ historical interactions and capture their interests. For instance, GRU4Rec (Hidasi et al., 2015) uses RNN, while Caser (Tang and Wang, 2018) uses CNN. SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) are based on the self-attention mechanism (Vaswani, 2017), which automatically learns the weight of each interaction.
In addition to innovations in model architecture, recent sequential recommendation models try to address the challenge of capturing rapidly changing user interests over time. For example, DROS (Yang et al., 2023a) introduces distributionally robust optimization (DRO) (Rahimian and Mehrotra, 2019) to enhance the model’s ability to handle out-of-distribution (OOD) problems caused by temporal changes. SURGE (Chang et al., 2021) uses GNNs to capture dynamic user interests within implicit interactions.
Furthermore, some studies have incorporated contrastive learning into sequential recommendations to alleviate the issue of sparse recommendation data. Such methods primarily focus on how to construct positive and negative samples within the contrastive loss. CL4SRec (Xie et al., 2022) constructs positive samples of the original sequence through sequence transformations such as item cropping, item masking, item reordering. DuoRec (Qiu et al., 2022) proposes a model-level data augmentation method based on dropout to achieve better semantic retention. DCRec (Yang et al., 2023b) introduces a contrastive learning loss based on conformity and interest disentanglement to address the popularity bias present in data augmentation. SoftCSR (Zhang et al., 2024c) extends contrastive learning to regional-level sample comparisons, offering more flexibility than single-sample contrasts, and employs adversarial contrastive loss to enhance model robustness.
The readers may refer to the excellent survey (Fang et al., 2020; Wang et al., 2019) for more details. However, the aforementioned methods do not consider the negative impact of noisy interactions on modeling the sequence, which is quite common in real-world scenarios.
@@ 5.2. Denoising Recommendation
In this section, we first review the current sequential recommendation denoising schemes, followed by a discussion of denoising recommendation under collaborative filtering scenarios.
5.2.1. Denoising sequential recommendation
@@
Regardless of whether explicit or implicit denoising methods are used, the most significant challenge stems from the lack of effective supervision signals from noise labels to train the denoising model. Existing methods often rely on heuristic designs inspired by prior knowledge to develop denoising algorithms, which requires substantial expert knowledge and lacks precision. Additionally, since these methods usually utilize supervision information from recommendation loss, they are still adversely affected by noise as the labels of the recommendation label (i.e., the next item) could also be contaminated. Our proposed method, LLM4DSR, addresses these issues by leveraging the extensive open knowledge of Large Language Models (LLMs) to achieve more effective denoising.
5.2.2. Denoising recommendation under collaborative filtering scenarios.
Considering other recommendation scenarios, most of work focuses on denoising under collaborative filtering scenarios. Due to the lack of labeled noise data, some works utilize prior knowledge about noise samples to aid in denoising for implicit feedback data. For instance, ADT (Wang et al., 2021a) and SGDL (Gao et al., 2022) utilize the phenomenon that models tend to fit clean samples more easily during the early stages of training, while noisy samples exhibit higher loss, to design denoising strategies. ADT simply sets a threshold to reduce the weights of samples with higher loss, whereas SGDL uses the information from the clean samples that the model has fitted to help train a sample weight adjustment model. DeCA (Wang et al., 2022) identifies noise by observing that different models exhibit significant prediction differences for noise samples. However, such prior knowledge is not always reliable. For example, Shu et al.(Shu et al., 2019) points out that samples with high loss are not necessarily noise and could also be hard samples. RGCF (Tian et al., 2022) and GraphDA (Fan et al., 2023) utilize node features extracted by GNNs to reconstruct the graph structure based on feature similarity, aiming to reduce the impact of noise signals in the graph. BOD (Wang et al., 2023a) employs a bi-level optimization approach to train a weight generator that automatically learns the weight of each sample. These methods often involve complex module designs and, due to the lack of labeled noise signals, their performance is still not guaranteed.
To the best of our knowledge, a recent study has also explored the use of LLMs for recommendation denoising in the collaborative filtering scenario (Song et al., 2024). However, our approach, LLM4DSR, significantly differs from LLMHD in several key aspects: (1) Different tasks: LLM4DSR addresses the task of sequential recommendation, while LLMHD focuses on collaborative filtering. Specifically, our work targets denoising sequential data, whereas LLMHD aims to differentiate between noisy samples and hard negative samples; (2) Different methods: LLM4DSR directly employs instruction tuning and uncertainty estimation to tackle the denoising task, while LLMHD directly uses untuned LLMs and crafts specific prompts to reassess samples, identifying those that deviate from traditional model predictions as noise. Moreover, LLMHD (Song et al., 2024) is a preprint available on arXiv as of September 16, 2024, and has not been formally published. Therefore, we consider our work to be parallel to LLMHD.
@@ 5.3. LLMs for Recommendation
The initial attempts directly use pre-trained LLMs as the backbone for recommendations to explore their zero-shot capabilities (Gao et al., 2023; Hou et al., 2024; Liu et al., 2024; Wang and Lim, 2023; Wang et al., 2024b). These methods perform not well due to the significant gap between recommendation tasks and the training tasks of LLMs. Subsequent research has organized recommendation data into prompt formats and then fine-tuned LLMs to enhance their recommendation performance, achieving better results (Bao et al., 2023a, b; Hu et al., 2024; Li et al., 2023a, 2024b, b, 2024a; Shi et al., 2024; Wang et al., 2024a; Zhang et al., 2023b, 2024b; Gao et al., 2024).
Besides generating recommendation results directly, LLMs can also serve as enhancers to improve the performance of conventional recommenders. Existing methods mainly employ LLMs as encoders to encode the semantic information of users and items into embeddings (Xi et al., 2023; Ren et al., 2024b; Wei et al., 2024; Wang et al., 2023b) or as an additional knowledge base (Yang et al., 2024; Qin et al., 2024). However, such methods largely fail to leverage the inherent reasoning capabilities of LLMs. To address this shortcoming, SLIM (Wang et al., 2024d) has enabled LLMs to generate Chain-of-Thought (CoT) data (Wei et al., 2022b) as additional input for downstream models. Since the content generated by LLMs is not optimized with recommendation task, this kind of information may not necessarily benefit traditional models. To address this, SeRALM (Ren et al., 2024a) considered optimizing the outputs of LLMs using recommendation loss by alignment training.
To the best of our knowledge, this paper is the first to explore the application of LLMs in the field of denoising sequential recommendation.